The wheeling tire company produces a batch of 5000 tires that includes exactly 200 that are defective.
a. if 4 tires are randomly selected for installation on a car, what is the probability that they are all good?
b. If 100 tires are randomly selected for shipment to an outlet, what is the probability that they are all good? Should this outlet plan to deal with defective tires returned by consumers?

Answer by stanbon(75887)

The wheeling tire company produces a batch of 5000 tires that includes exactly 200 that are defective.
a. if 4 tires are randomly selected for installation on a car, what is the probability that they are all good?
-------
P(defective) = 200/5000 = 1/25 = 0.04
P(good) = 0.96
---------------------
P(4 good) = 0.96^4 = 0.8493
=====================================
b. If 100 tires are randomly selected for shipment to an outlet, what is the probability that they are all good?
P(100 good) = 0.96^100 = 0.0169
-------------------------------------
Should this outlet plan to deal with defective tires returned by consumers?
I'll leave that to you.
Cheers,
Stan H.
============